(−4, − 7𝜋/6)
Find three additional polar representations of the point, using −2𝜋 < 𝜃 < 2𝜋.
(r, 𝜃) =
(r, 𝜃) =
(r, 𝜃) =

Answer by greenestamps(13215)



(1) Keep the "-4" and add or subtract any even multiple of pi (i.e., a multiple of 2pi) to the given angle;

OR

(2) Change the "-4" to "4" and add or subtract any odd multiple of pi to the given angle.

For example....

(-4, -7pi/6+4pi) = (-4, 17pi/6)
or
(4,-7pi/6-9pi) = (4,-61pi/6)

But those two do not use angles in the specified range of -2pi to +2pi.